OpenStudy (ahristudent):

Hi, I need help on this math question that's going to be on my exam... in about an hour. I need help! I give medals! ** This is the question: Admission to a basketball game was $1 for students and $2 for adults. The number of tickets sold for students was 80 more than the number of adult tickets sold. If the total amount of money came out to $980, how many of each ticket was sold? It's very hard for me as I don't know how to set this up and actually do it, I'm not the best in algebra, I tried but got weird answers. Thank you !

OpenStudy (dayakar):

let the no. of students =x no.of adults = y admission fee for 1 student =1$ admission fee for 1 adult =2$ total amt = 980 x + 2y =980 ------(1) students = adults +80 [students are 80 more than adults] x= y+80 ------(2) substitute (2) in (1) y+80 +2y = 980 3y +80 =980 3y = 980-80 3y = 900 y= 900/3 y = 300 there for no of adults are y = 300 plug y= 300 in (2) you can get x

OpenStudy (ahristudent):

Thank you so much!! <3